OpenStudy (anonymous):

really quick help for fan/medal? Write the product as a trinomial (k – 1)(k – 4) Is it k^2 – 5k – 4

OpenStudy (whpalmer4):

No. It isn't. What is -1 * -4

OpenStudy (anonymous):

(k – 1)(k – 4) =k^2-4k-1k+4 =k^2-5k+4 (:

OpenStudy (anonymous):

Ah okay thanks so much! :D